OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (johnweldon1993):

So since we know that 'y' = the total number of votes cast....and we are told that that is 125000 people...we can plug that in this equation and solve for 'x' which is the number of votes she received... \[\large x - 62,500 = 0.25y\] after replacing y with 125000 we have: \[\large x - 62,500 = 0.25(125,000)\] Now simplify the right hand side... \[\large x - 62,500 = 31250\] And now just add 62500 to both sides of the equation... \[\large x = \space ?\]

OpenStudy (anonymous):

93,750?

OpenStudy (johnweldon1993):

That is what I get! :)

OpenStudy (anonymous):

thxs! hey can you help me with one more?

OpenStudy (johnweldon1993):

That new question you just posted? sure I was going to go there now anyways lol :)

OpenStudy (anonymous):

thxs! :)